OpenStudy (anonymous):

(i) Using the expansions of cos(3x − x) and cos(3x + x), prove that 1/2(cos 2x − cos 4x) ≡ sin 3x sin x.

OpenStudy (hoblos):

do you know the expansions of cos(3x − x) and cos(3x + x) ?

OpenStudy (anonymous):

sum/difference formula for cosine...

OpenStudy (hoblos):

cos(a-b) = (cosa)(cosb) + (sina)(sinb)

OpenStudy (hoblos):

cos(a+b) = (cosa)(cosb) - (sina)(sinb)

OpenStudy (hoblos):

try using these formulas here...

OpenStudy (callisto):

cos(3x − x) = cos3xcosx + sin3xsinx - cos(3x + x) =cos3xcosx - sin3xsinx ----------------------------------- cos 2x-cos 4x =sin3xsinx-(-sin3xsinx) cos 2x - cos4x = 2sin3xsinx 1/2 (cos 2x - cos4x) = sin3xsinx

OpenStudy (callisto):

It's just the concept only, write a better proof :)

OpenStudy (hoblos):

yup.. thats what you had to do or simply replace cos 2x by cos3xcosx + sin3xsinx and cos 4x by cos3xcosx - sin3xsinx and it can be proved...

OpenStudy (anonymous):

I'll look at this soon. I need to do something... Thanks for helping!! I'm getting some of it

OpenStudy (callisto):

LS = 1/2(cos 2x − cos 4x) = (1/2) [ cos(3x − x) - cos(3x + x)] = (1/2) [ (cos3xcosx + sin3xsinx) - (cos3xcosx - sin3xsinx)] = (1/2) (2sin3xsinx) = sin3xsinx = RS Therefore, 1/2(cos 2x − cos 4x) ≡ sin 3x sin x